OpenStudy (anonymous):

Let T be an m-cycle ( 1 2 3 ... m ) , show that T^i is an m-cycle iff m,i are relativity prime

OpenStudy (anonymous):

if there is more than a prove , plz , add it

OpenStudy (cruffo):

I think I have part of the proof: I started with a few small cases, for example, suppose m=5. T=(12345) T^2=(13524) T^3=(14253) T^4=(15432) Let T=(123⋯m) be an m-cycle Notice that for each 1≤k≤m, T^i: k→(k+i) mod m Suppose T^i is an m-cycle, then T^i=(1 (1+i)mod m (1+2i)mod m ⋯ (1+(m−1)i) mod m) Furthermore, for 1≤n≤m−1, 1+ni≢1(mod m) Thus ni≢0(mod m) So m and i are relatively prime.